OpenStudy (anonymous):

integrate 1/(x^3(sqrt(x^2-1))

OpenStudy (anonymous):

here is the deal you com and help me on mkine and I will help u on yrs deal or no deal/

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ x ^{3}(\sqrt{x ^{2}-1}) }\]

OpenStudy (anonymous):

i help with wat?

OpenStudy (solomonzelman):

substitute. x=sec(u) dx = tan(u)sec(u) du and then, √(x^2-1)=√(sec^2u-1)=√tan^2u = tan u u= sec^(-1)x

OpenStudy (anonymous):

ok let me see. I did get somewhere , let me write where i got to

OpenStudy (solomonzelman):

sure

OpenStudy (solomonzelman):

(I wondered if I can help with integration before I actually even learn derivatives in class)

OpenStudy (solomonzelman):

you are patient :) writing it out. But just saying, that no matter what you get, derive it to check.

OpenStudy (anonymous):

\[x =\sec theta\] \[dx=\sec theta tan theta d theta\] \[\int\limits_{}^{}\frac{ \sec \theta \tan \theta }{\sec ^{3}\theta (\sqrt{\tan ^{2}\theta}) }\] \[\int\limits_{}{\frac{ 1 }{ \sec ^{2}\theta }}\] \[\int\limits_{}^{}\cos ^{2}\theta\] \[= (1/2 )(\theta + \sin \theta \cos \theta)\] \[\frac{ 1 }{ 2}\sec ^{-1}x+ what?\]

OpenStudy (anonymous):

i dont know how to make 2sin@cos@ to be\[\frac{ \sqrt{x ^{2}-1} }{ 2x }\]

OpenStudy (anonymous):

sorry its \[\frac{ 1 }{ 2 }\sin \theta \cos \theta\]

OpenStudy (solomonzelman):

well cos^2u is correct. then re-write cos^2u as 1/2cos(2u)+1/2

OpenStudy (anonymous):

i dont know how to make it be in fraction form, if my working is alryt. I have cut some stages short

OpenStudy (anonymous):

thats what i did for the integral and i got\[1/2(\theta +\sin \theta \cos \theta)\]

OpenStudy (anonymous):

\[x=\sec \theta\] \[\theta=\sec ^{-1}x\] \[then \frac{ 1 }{ 2 }\theta = makes \theta =2\sec ^{-1}x\]

OpenStudy (solomonzelman):

I am getting S 1/2cos(2u) +1/2 du S 1/2cos(2u) du + S 1/2 du sin(2u)/4 + u/2

OpenStudy (solomonzelman):

and final answer then is 1/2 (x^2) / √(x^2-1) - (1/2) tan^(-1) (1/ √(x^2-1)) + C

OpenStudy (anonymous):

uhmmm, let me retrace your steps. and pliz just hint me on the following \[\int\limits_{}^{}\frac{ 1 }{ 1+2e ^{x}+e ^{-x} }\]. What type of substitution or trgi should I make use of? Partial fractions or some factorisation?

OpenStudy (solomonzelman):

this is what I would do. u=e^x

OpenStudy (anonymous):

ok, let me c

OpenStudy (solomonzelman):

I think I have to go right now -:( Well... don't worry about me, I am not such a great helper anyways... bye

OpenStudy (anonymous):

thanx man. Just that the u substitution u opted for is a dark tunnel to me. i dont see the light